OpenStudy (anonymous):

What is the missing side length of an irregular quadrilateral with side lengths 21.7 centimeters, 34.6 centimeters and 29.2 centimeters, if the perimeter is equal to the perimeter of a regular hexagon with side length 18 centimeters?

hartnn (hartnn):

perimeter of a regular hexagon is 6*side....so now u know perimeter of irregular quad...with only one side unknown....quite simple :)

OpenStudy (anonymous):

perimeter of hexagon of side 18 = 6*18 = 108

OpenStudy (anonymous):

21.7 + 34.6+ 29.2 + x = 108 find x?

OpenStudy (anonymous):

thx so much for your help! :D

OpenStudy (anonymous):

welcome...)_